state and proove bernallis theorem...........????????????????

Grade:11

1 Answers

Peddireddy Hari Kumar
8 years ago
Bernoulli’s equation states that the work done on a unit volume of
fluid by the surrounding fluid is equal to the sum of the changes in kinetic and
potential energies per unit volume that occur during the flow.
dW = p 1 A 1 ds 1 - p 2 A 2 ds 2 = (p 1 - p 2 )dV
The net change in kinetic energy dK during time dt is
dK = ½ ∂ dV(V 22 - v 12 )
dU = ∂ dV g(y 2 - y 1 )
dW =
dK + dU, we obtain
(p 1 - p 2 )dV = ½ ∂ dV(V22 – V 12 ) + ∂   dV g(y 2 - y 1 )
p 1 - p 2 = ½ ∂ (v22 - v12 ) + ∂ g(y2 - y1 )
p1 +∂gy1 + ½ p v12 + ∂gy1 + ½ ∂v12 = p2 + ∂gy2 + ½ ∂v22
(Bernoulli’s equation)
{∂= density}
thanq...